Montecino Corporation uses two different types of labor to manufacture its product. The types of labor, Cutting and Setup, have the following standards:

Labor Type Standard Mix Standard Unit Price Standard Cost
Cutting 500 hours $7.50 per unit $3,750
Setup 125 hours 5.00 per unit $ 625
Yield 3,125 units


During July, the following actual production information was provided:

Labor Type Actual Mix
Cutting 4,375 hours
Setup 1,875 hours
Yield 28,125 units


What is the Labor mix variance?

a.$1,563(U)

b.$3,750(F)

c.$3,750(U)

d.$1,563(F)
Answer #1

Actual Hours = 4375 Hours + 1875 Hours= 6,250 Hours

Standard Hours = 500 Hours + 125 Hours = 625 Hours

Actual Hours at Standard Mix

Cutting = 6,250 x (500/625) =5,000

Set up = 6,250 x (125/625) = 1,250

Labor Mix Variance = (Actual Hours used at Standard Mix - Actual Hours used at Actual Mix ) x Standard Hourly rate

Cutting = (5,000 - 4,375) x 7.50 = 4,688 (Favorable)

Set up = (1,250 - 1,875) x 5 = 3,125 (Unfavorable)

Total Labor mix Variance = 1,563 (Favorable)

So Answer is d.$1,563(F)